OpenStudy (anonymous):

what is the measure of one angle in a regular 30-gon

OpenStudy (anonymous):

(30-2)*180 / 30

OpenStudy (anonymous):

Sum of interior angle = (n-2)*180 one angle = sum / no. of angle